OpenStudy (anonymous):

the perimeter of a rectangle is 26 ft. The length is 7 ft longer than the width. find the dimensions. write the system of linear wquations ans solve the resulting system 2x+2y= write the first equation

OpenStudy (anonymous):

2x+2y = 26 x+7=y 2x+2(x+7) = 26 2x+2x+14=26 4x=12 x=3 y=10

OpenStudy (anonymous):